RC Update and Run plan for Sep 13,2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are M-F at 4:15 PM

To join the Meeting:  https://bluejeans.com/861439920
To join via phone :  +1.408.740.7256 (US (San Jose)), +1.888.240.2560 (US Toll Free))

Sunday 13 Sept 2020.

General updates

(See d2n Run Plan 2020 for broader status/overview.)

  • Taken about 12 hours of data in last 24; Accelerator struggling with a series of hardware problems
  • Down from 1am until 4pm for a series of issues with the chopper in the Injector [1] and south linac [2].
    • Problem with 1L17 that took us down yesterday doesn't look fixed either [3] (MCC leaving it alone for now).
    • Opportunistic access Sun morning to reset spectrometer rotation (tripped on Friday around the same time as the HMS dipole)
  • Set spectrometers to Kin-A, Kin-X Sunday morning and will continue with 90-deg running when beam returns [
  • Max Target Tommy polarization @ 90 since Sept 10 was reached this morning (after being down for a shift) at 38.8% [4]
  • Junhao reports the spectrum analyzer parts are ready to go for Monday
    • RadCon has been looped in and will support us
  • Target should be rotated to 270 deg on Monday

Notes for the Future

  • Bill, Murchhana check compass/field measurements and will report back at Sun RC meeting
  • Melanie, Murchhana please check good electron yields per Coulomb for Kin-A/X runs with T1 vs. T3
  • Beam Studies planned for Monday @ 10am to ~9pm (including recovery), Sept 14; with opportunity for 5-pass running mid-day
    • 10am: Hall C will enter Controlled Access as planned
    • Target work (2 tasks currently planned; RadCon support needed)
      • Replace a spectrum analyzer mounting plate
      • Take a compass measurement of the field at the carbon foil ladder position if possible [TBD]
      • Rotate target to Transverse/270
    • SOLID cherenkov config change (in parallel)
    • Lock-up Hall when done and take data (Ref cell, Optics?) until 2nd part of beam studies begin
    • 4:15pm - 9:15pm: Beam studies, part 2
    • Requires beam to Hall C dump, but no production running will be practical (as I understand it)
    • 9:15: Continue production running in Kin-A, Kin-X at transverse/270 when beam returns

Pending Issues

  • Check EPR calibrations (particularly for 270deg setting, but others too) [Melanie]
  • Check field direction measurements (Compass and field probe) [Murchhana, Bill]
  • Magnet EPICS PV timeout issue; perhaps reboot skylla10 on Monday?
    • Some Magnets Channels Lost Connection Briefly [5] (Saturday, 2pm)
    • Also had to add more error checking to the angle reporting in the CODA run-start GUI last week too (same softIOC as above)
    • 'Heartbeat' issue reported by Steve Lassiter last week
  • Target polarization for Tommy @ 270 is low (mid to high 30s); Tommy @ 90 seems to have plateaued at 38%
    • Spectrum analyzer collimator was broken at some point and present optics setup may be allowing transient light to interfere with spectrum measurement and QWP scans
    • JP wonders if the field direction is off for the transverse settings (Bill, Murchhana: see note in Target group chat from late Thursday)
  • Beam energy drifts
    • Just something to keep an eye on. Has not been a problem lately.
    • Hall C 5-pass polarization goes down as energy (HALLC:p) goes up [6]

Opportunistic Jobs/Access

Current Run Plan

Sunday DAY → Monday OWL

  • Sunday at 8am Move spectrometers to Kin-A, Kin-X and continue Transverse/90 running there [Done]
    • Update correction coil settings for Kin-X HB field [Done]
    • NOTE: Take 1 run with T1(3/4) trigger; then 3 runs using T3(EL_clean) trigger; Then repeat that cycle.
      • Be sure to monitor and set the prescales correctly.
      • Put T1 or T3 in the CODA comments field for the runs.
      • Note the T1 or T3 trigger type for each run in the Shift Summary and Run database using the add_run_comment command.

Monday DAY

  • 10am: Hall C will enter Controlled Access as planned
    • Target work (2 tasks currently planned; RadCon support needed)
      • Replace a spectrum analyzer mounting plate
      • Take a compass measurement of the field at the carbon foil ladder position if possible [TBD]
      • Rotate target to Transverse/270
    • SOLID cherenkov config change (in parallel)
  • Lock-up Hall when done and take data (Ref cell, Optics?) until 2nd part of beam studies begin
  • 4:15pm - 9:15pm: Beam studies, part 2
    • Requires beam to Hall C dump, but no production running will be practical (as I understand it)
  • 9:15: Continue production running in Kin-A, Kin-X at transverse/270' when beam returns
    • NOTE: Take 1 run with T1(3/4) trigger; then 3 runs using T3(EL_clean) trigger; Then repeat that cycle.
      • Be sure to monitor and set the prescales correctly.
      • Put T1 or T3 in the CODA comments field for the runs.
      • Note the T1 or T3 trigger type for each run in the Shift Summary and Run database using the add_run_comment command.

Monday SWING → Tuesday OWL

  • Continue production data in Kin-A, Kin-X with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs.
    • NOTE: Take 1 run with T1(3/4) trigger; then 3 runs using T3(EL_clean) trigger; Then repeat that cycle.
      • Be sure to monitor and set the prescales correctly.
      • Put T1 or T3 in the CODA comments field for the runs.
      • Note the T1 or T3 trigger type for each run in the Shift Summary and Run database using the add_run_comment command.

Tuesday DAY → Tuesday SWING

  • Continue production data in Kin-A, Kin-X with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs.
    • NOTE: Take 1 run with T1(3/4) trigger; then 3 runs using T3(EL_clean) trigger; Then repeat that cycle.
      • Be sure to monitor and set the prescales correctly.
      • Put T1 or T3 in the CODA comments field for the runs.
      • Note the T1 or T3 trigger type for each run in the Shift Summary and Run database using the add_run_comment command.
-- General plan below this point is fixed, but details may be adjusted as beam/target requires --

Wednesday OWL → Wednesday SWING

  • Moller measurement on Wednesday DAY
    • This will also be used to infer polarization at 1-pass (if it fails, we'll try again on Thursday DAY)
  • Wednesday OWL or DAY during Moller Move spectrometers to Kin-C, Kin-Z and continue Transverse/270 running there
  • Continue production data in Kin-C, Kin-Z with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs when Moller complete.
    • All runs for these settings use the T1 trigger only

Wednesday OWL → Wednesday SWING

  • Continue production data in Kin-C, Kin-Z with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs.
    • All runs for these settings use the T1 trigger only

Thursday OWL → Friday OWL

  • Continue production data in Kin-C, Kin-Z with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs.
    • All runs for these settings use the T1 trigger only
  • 6 am Move to Controlled Access (w/ RadCon support) to begin transition to 1-pass

Friday DAY

  • (Started at 6am) Open target enclosure
    • Install collimators (JP, RadCon)
    • Check on EPR photo-diode status, likely replace it? (JP, RadCon)
  • Ensure fast raster is reconfigured correctly (Expert: Jones, Gunning)
  • Rotate to minimum angles for HMS (11.7) and SHMS (8.5) (S. Lassiter, A. Comer)
  • Configure target for Longitudinal/0 or Longitudinal/180 (TBD)
  • Close up and wait for MCC to establish (1-pass) beam to the Hall

Friday SWING → Monday

  • See task list at Run Plan for 1-pass for more details
    • Initial priority will be PbPt polarization cross check with longitudinal pol. elastics.
    • These have lower priority but will likely be pursued while target spins up
      • 3He, N2 density measurements
      • H2 xsec measurement
      • Pressure curves


Standing Orders

  • Change the IHWP at least every 8 hours when running on glass cells. Call Halls A and B in advance. If there are any issues, call the RC
  • Ensure that all information in the CODA run-start GUIs is correct for every run!
  • If beam goes away for more than 3h, or MCC does operations on the beam that may change the beam tune, then move target to No Target position,
    • verify Harps scans at 3H07A and 3H07B are within spec, and
    • ensure beam is stable (with all locks ON) at 0, 0mm before continuing.
    • Beam recovery procedure has more detailed procedure if needed.
  • New beam positions have been verified and established to be (0,0) mm for all targets.
    • Monitor that beam is at (0,0) using the stripchart on the Wall screen and verify no scraping on the glass cells using the Analysis monitor.
  • Check the position locks are on.
  • Check that CODA is running normally (had one strange hangup lately)
  • Check online replays shortly after starting runs, & full replay results too.
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Follow the updated guidelines in the How-to#Spectrometer_Magnets Spectrometer Magnets How to
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • If any magnet trips, snapshot the associated power supply window and post it to hclog before resetting the interlock.
  • Monitor the beam position on target and call MCC immediately if you see large excursions (1 mm level).
  • When rotating the spectrometer, use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instructions are here: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT

Useful Links

PolHe3 Polarization Measurements

Misc Links